Actions
Ops/Comm Scheduling meeting¶
Location : CR11 in R3 @ 15:30 pm on Tuesday 10th
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Agenda¶
- Actions from last meeting
- Generate appropriate table to store absorber pressure and material information (Janusz) : (almost done)
- DS readback current requires calibration (DL) : stands until close to powerup
- Issues arising from last data taking run need to be identified and actioned (Boyd) : Stands (but I have a short list)
- Investigate the state of the refurbished air end to see what we get for our money (Vicky)
- Colin will put the ISIS dates for run-ups and machine work into his schedules to better reflect available access time. (Done)
- Some form of water system schedule will be made available (Colin)
- Henry, Craig, Paul and Pierrick to look into the water detectors in the trench
- DC report
- Work plan for the next week : http://micewww.pp.rl.ac.uk/attachments/6185/2016-05-03-MICE-Hall-Plan.xlsx
- QP/QD work : http://micewww.pp.rl.ac.uk/attachments/6181/MICE_Electrical_Work_List_25APR16.docx
- Water system work
- FC cooldown
- Bellows re-installation
- AOB
- DONM : 17/3/2016 IN CR05 R18
Minutes¶
- *Present : * Paul H, Steve B, Mark T, John G, Henry N, Ed O, Josef F
- Actions from last meeting
- Generate appropriate table to store absorber pressure and material information (Janusz) : (almost done)
- DS readback current requires calibration (DL) : stands until close to powerup
- Issues arising from last data taking run need to be identified and actioned (Boyd) : Stands (but I have a short list)
- Investigate the state of the refurbished air end to see what we get for our money (Vicky)
- Colin will put the ISIS dates for run-ups and machine work into his schedules to better reflect available access time. (Done)
- Some form of water system schedule will be made available (Colin)
- Henry, Craig, Paul and Pierrick to look into the water detectors in the trench : Still being worked on. Stands
- DC report :
- Work on Linde plant done successfully and DS fridge cooling down
- Work on water continues.
- Daresbury assisted in testing QP into short last week. This was successful.
- Target has an anomalous light loss which Paul and the target team is investigating. The target team will be here in
the first week in June to investigate this issue.
- Work plan for the next week : http://micewww.pp.rl.ac.uk/attachments/6185/2016-05-03-MICE-Hall-Plan.xlsx
- Action Item Check ISIS spreadsheet (Boyd)
- QP/QD work : http://micewww.pp.rl.ac.uk/attachments/6181/MICE_Electrical_Work_List_25APR16.docx
- Resistor dump packs being installed this week
- Attempting to get everything commissioned for sign-off by the week of the 16th May
- A soak test of the system will be done tomorrow night (night of the 11th)
- Water system work
- Brackets are in and the pipework from the chiller into the Hall is being installed
- Chiller can't be positioned until the road is fixed
- FC cooldown
- Bellows are going in and FC has been pumped down
- Prepare for cooling down later this week subject to availability of water cooling system
- Must ensure that load cells and temperature sensors are all connected up.
- Bellows re-installation
- See above
- AOB
- Synch will be closed on 9/10th at end of shutdown period. Need to change of target timing at some point. Should organise
with the ISIS to do this in the same week as we do target work. Paul will take care of this. - HPC are coming to work on the compressor repair : on 7/8/9 August -> work plan.
- Synch will be closed on 9/10th at end of shutdown period. Need to change of target timing at some point. Should organise
Updated by Boyd, Steven about 7 years ago ยท 5 revisions